OpenStudy (anonymous):

what is the inverse of f(x)=-2x+7

OpenStudy (anonymous):

x-7/-2=f-1(x)

OpenStudy (cwrw238):

y = -2x + 7 x = 7 - y ---- 2 f(-1)x = (7-x) / 2

OpenStudy (anonymous):

2 or -2?

OpenStudy (cwrw238):

both are correct (7-x) / 2 is same as (x-7)/-2 ( i dont like negatives!)

OpenStudy (anonymous):

f(x)=-2x+7 2x = -y + 7 ->x = (-y + 7)/2 => f^-1 (x) = -x/2 + 7